Say that f(x) = 2x - x. What would be the result of f(5)


Answer by fcabanski(1391)

f(5) = 2*5 - 5 = 10-5 = 5.


You can also write f(x) as x, since 2x - x = x.


Thus f(x) = x so f(5) = 5.